OpenStudy (anonymous):

The displacement (in centimeters) of a particle moving back and forth along a straight line is given by the equation of motion s = 2 sin πt + 2 cos πt, where t is measured in seconds. (Round your answers to two decimal places.) (a) Find the average velocity during each time period. (i) [1, 2]

OpenStudy (dumbcow):

avg velocity: \[\frac{1}{b-a} \int\limits_a ^{b} s'(t) dt\]

OpenStudy (dumbcow):

\[= \frac{1}{b-a} [s(b) -s(a)]\]

OpenStudy (dumbcow):

essentially its the slope formula, so for interval [1,2] \[\frac{s(2) -s(1)}{2-1}\]